OpenStudy (anonymous):

A rectangle has the perimeter of 26 inches. The length is 3 inches longer than the width. Find the dimensions of the rectangle A. 5 in by 8 in B. 6 in by 9 in C. 2 in by 5 in D. 8 in by 11 in Please explain how to solve

OpenStudy (anonymous):

Do you know the formula for perimeter of rectangle?

OpenStudy (anonymous):

Do you know the formula for perimeter of rectangle?\[\large \color{green}{Perimeter_{rectangle} = 2 (Length + Width)}\]

OpenStudy (anonymous):

I expected a reply..

OpenStudy (anonymous):

Length = 3 + Width.. Or \(l = 3 + w\) \(26 = 2(l + w) \implies 26 = 2(3 + w + w) \implies 26 = 2(3 + 2w)\)

OpenStudy (anonymous):

Divide by \(2\) both the sides, then subtract \(3\) from both the sides and then divide by \(2\) both the sides...